OpenStudy (shadowcain):

The product of 9 and a number is 12 less than three times that number. What is the number?

Directrix (directrix):

Let x be the number. The product of 9 and the number is: 9x

Directrix (directrix):

Three times that number is 3x 12 less than three times that number is 3x - 12

OpenStudy (shadowcain):

9x=3x-12 6x=-12 x=-2

Directrix (directrix):

Your task: To solve the equation: 9x = 3x - 12 @ShadowCain

OpenStudy (shadowcain):

9x = 3x - 12 = -2

Directrix (directrix):

>> 9x = 3x - 12 = -2 You lost the variable x.